a) find and classify the critical points of the functions f(x) = 2x^3 + 3x^2 - 12 x +1 into maximum, minimum and inflection points as appreciate.

(b) The sum of two positive numbers is S. find the maximum value of their product.



1
Expert's answer
2021-07-25T10:59:27-0400

"i)\ 

f(x)=2x^{3}+3x^{2}-12x+1"


For this "f(x)" , we have to find critical points , maximum and minimum points and inflections also of this given functions .


"f^{'}(x)=6x^{2}" "+\ 6x-12"


To find the critical values , we differentiate and find values of "x," that is we put "f^{'}(x)=0" ,



"f^{'}(x)<0\implies" "f(x)" is decreasing


"f^{'}(x)=0\implies" "f(x)" is stationary


"f^{'}(x)>0\implies" "f(x)" is increasing.


"f^{'}(x)=6x^{2}+6x-12"


"f{'}(x)=0"


"\implies" "6x^{2}+6x-12=0"


At a critical point , "f{'}(x)=0"


"\implies" "x^{2}+x-2=0"


"{\therefore}" "x=-2,1"


"x=-2\implies" "f(-2)=20"


"x=1\implies" "f(1)=-7"


Differentiating , with respect to "x,"


"f^{''}(x)=12x+6"


"x=-2\implies f^{''}(-2)=" "-8<0" , maximum



"x=1\implies\ f^{''}(1)=18>0" , minimum



Inflection points , for the function -


"f(x)=2x^{3}+3x^{2}-12x+1"


An inflection point is a point on a curve where concavity changes from concave up to concave down or vice versa .


"f^{'}(x)=6x^{2}+6x-12"


"f^{''}(x)=12x+6"


Inflection point also has same meaning as where "f^{''}(x)" switches sign -


"=12x+6=0"


"=x=\dfrac{-1}{2}"

Putting the obtained "x" in the equation "f(x) ," we get -


"y=\dfrac{15}{2}"


So point of inflection are -"(\dfrac{-1}{2},\dfrac{15}{2})"


"ii)" Let the two numbers be "x\ and \ y." .


"=x+y=S" .................................1)


"=" "y=S-x"

their product is given as -


"=xy=x(S-x)=Sx-x^{2}"


"f(x)=Sx-x^{2}"


"f^{'}(x)=S-" "2x"


putting "f^{'}(x)=0" , we get -


"x=\dfrac{S}{2}"


Putting the value of "x\" in equation 1) , we get ,


"y=\dfrac{S}{2}"


The maximum value of product can be written as -


"xy" , Putting the value of "x\ and \ y-"


"=\dfrac{S}{2}\times\dfrac{S}{2}=\dfrac{S^{2}}{4}"


Which is required product